# The InSAR absolute phase amid singularities

### v1

@@ 6.3 Loops in the punctured complex plane
[…]
Evaluation is facilitated by $d\varphi_{\mathbb{C}^{+}}$ being closed [21], $\mathrm{d}d\varphi_{\mathbb{C}^{+}}=0$. (From this, the earlier claim that $d\varphi_{W^{+}}$ is closed follows, as the pullback commutes with the exterior derivative.) The one-form $d\varphi_{\mathbb{C}^{+}}$ in $\mathbb{C}^{+}$ is the canonical example of a closed one-form that is not exact, i.e., it cannot be written as the exterior derivative of a zero-form [8].
[…]
The winding number $N(h)\in\mathbb{Z}$ counts the net number of turns $h$ makes around the origin and is positive for counterclockwise turns. It follows that the unwrapped phases along two paths differ by an integer multiple of $2\pi$, namely $N(\gamma\circ g)$ of the corresponding loop $g\in W^{+}$.
Two loops in $\mathbb{C}^{+}$ have the same winding number if and only if they are homotopic to each other [8], with the basepoint fixed. Consequently, a loop with winding number $N$ and basepoint $1+0\mathsf{j}$ can be deformed into the circular loop $h_{N}(s)\equiv\cos(2\pi Ns)+\mathsf{j}\sin(2\pi Ns)$ with $s\in[0,1]$. Thus, integrating $d\varphi_{\mathbb{C}^{+}}$ around a loop $h$ gives zero if and only if $N(h)=0$, or equivalently, if and only if $h$ can be shrunk continuously to the constant loop $h_{0}$, i.e. $h\simeq h_{0}$.
@@ 6.4 Identifying patchability in $W^{+}$
The absolute phase singularities $A^{\prime}$ are thus patchable if any arbitrary loop $g\in W^{+}$ with basepoint $\mathbf{w}_{1}$ gets mapped to a loop $\gamma\circ g$ in the $\mathbb{C}^{+}$ with $N(\gamma\circ g)=0$. No loop $g\in W^{+}$ can have $N\gamma\circ g)\neq 0$. This raises two questions, namely what characterizes loops $g$ in $W^{+}$ with $N(\gamma\circ g)\neq 0$, and how can we tell whether such loops exist?
Two properties characterize a loop $g\in W^{+}$ with $N(\gamma\circ g))\neq 0$. First, it encircles a phase singularity (or other hole in $W^{+}$) such that it cannot be continuously shrunk to the constant loop $\varepsilon_{\mathbf{w}_{1}}$. If it were homotopic to $\varepsilon_{\mathbf{w}_{1}}$, the integral would evaluate to zero due to (26), which in turn implies $N(\gamma\circ g)=0$ due to (27), a contradiction. Second, it intersects $A^{\prime}$, the set of absolute phase singularities on which the phase is defined. To see this, assume it did not, so that its support is in $W^{\mathrm{a}}$. However, $d\varphi_{W^{+}}=\mathrm{d}\phi^{\mathrm{a}}$ and thus exact in $W^{\mathrm{a}}$, implying $N(\gamma\circ g)$ = 0, yielding a contradiction.
[…]
Whether absolute phase singularities $A^{\prime}$ can be patched thus relates to the topology of the entire space $W^{+}$, as described using homotopy through $\pi_{1}(W^{+},\mathbf{w}_{1})$. Conversely, the absolute phase itself and consequently whether a point $\mathbf{w}_{2}$ is an absolute phase singularity, $\mathbf{w}_{2}\in A^{\prime}$, only depends locally on the absolute path.
@@ 7 Reconstructing the absolute phase
Under what conditions can the absolute phase be reconstructed from discrete InSAR measurements that do not allow for its direct observation? A general two-step approach [6, 2] for a multidimensional interferogram is to i) unwrap the phase to obtain the absolute phase up to a uniform offset, an integer multiple of $2\pi$; and ii) estimate this offset using such strategies as multifrequency observations.
@@ 7.1 Multidimensional interferograms
In a multidimensional interferogram, the absolute phase may be reconstructed through the slow variation of the phase along these dimensions, most commonly space or time [7]. In an ideal scenario, phase unwrapping reconstructs the absolute phase up to an integer multiple of $2\pi$, the estimation of which is deferred to Sec. 7.2. For now, the focus is on the conditions under which the absolute phase can be reconstructed up to a single constant by unwrapping, treating idealized continuous and then discretely sampled observations.
@@ 7.1.1 Continuous unwrapping amid singularities
The complex coherence $\gamma$ is assumed to vary smoothly in the space $X\cong S\subset\mathbb{R}^{n}$. For instance, $n=2$ for a spatial interferogram. Discontinuities in $\gamma$ (e.g., layover) are assumed to have been excised from $X$. The set $X^{+}=X\setminus\gamma^{-1}\left[\{0\}\right]$, obtained by removing phase singularities, is open, path connected and locally path connected [10]. The wrapped phase $\varphi=\mathrm{arg}\gamma$ is assumed to arise congruently from a smooth absolute phase field $\phi^{\mathrm{a}}:X^{\mathrm{a}}\rightarrow\mathbb{R}$, where $X^{\mathrm{a}}=X\setminus A$, the space $X$ excluding all absolute phase singularities $A$. This means that $\gamma=\rho\exp\left(\mathsf{j}\phi^{\mathrm{a}}\right)$ on $X^{\mathrm{a}}$ with smooth $\rho$. For instance, in differential InSAR with given observation times $t_{1}$ and $t_{2}$, a smooth $\phi^{\mathrm{a}}(\mathbf{x})$ over $X$ is obtained whenever the unobserved coherence $\gamma(\mathbf{x},t)$ at $t\in[t_{1},t_{2}]$ with fixed primary $t_{1}$ is a smooth function of space $\mathbf{x}$ and time $t$; however, vanishing coherence at $\mathbf{x}$ and some intermediate $t$ induces an absolute phase singularity at $\mathbf{x}$ at secondary time $t_{2}$.
@@
[…]
@@
The absolute phase can be reconstructed up to a uniform offset if all unwrapping paths are restricted to $X^{\mathrm{a}}$. Reconstructability in a path connected space $X^{\mathrm{a}}$ follows because the observed differential form $d\varphi_{X^{+}}$ is identical to $\mathrm{d}\phi^{\mathrm{a}}$ on $X^{\mathrm{a}}$ due to the assumed smoothness and congruence. The goal, then, is to avoid absolute phase singularities during unwrapping.
@@
Are absolute phase singularities related to branch cuts? Branch cuts are barriers erected during phase unwrapping to circumvent path dependence [27, 2]. For instance, in Fig. 8c, the two phase singularities could be connected by a branch cut, which enables recovery of an unwrapped phase field outside the branch cut. However, this unwrapped phase bears no resemblance to the absolute phase in Fig. 8d. To recover the absolute phase, branch cuts should be placed at absolute phase singularities, but the examples in Fig. 8 show how difficult than can be based on observations. Conversely, the strategy of linking the two phase singularities with a branch cut is appropriate when path dependence is induced by localized noise, to be studied in discrete data.
@@ 7.1.2 Discrete unwrapping amid singularities
Path-based phase unwrapping proceeds along a series of segments coordinate-parallel segments connecting adjacent lattice points. A space $X_{\mathrm{d}}\subset\mathbb{Z}^{n}$ is called lattice path connected if any pair of points can be connected by such a path traversing adjacent pixels. Phase unwrapping along a unit step in $x_{i}$, denoted by $\Delta_{i}$, evaluates
@@
known as Itoh’s criterion [28]. Itoh’s criterion requires the absolute phase to change slowly between adjacent lattice points and can be violated due to steep gradients (undersampling) or apparent discontinuities [6]. The required slow variation warrants re-interpretation amid absolute phase singularities.
Itoh’s criterion can be violated due to absolute phase singularities, in addition to rapid phase changes or noise. An example is shown in Fig. 8a. Assuming the continuous coherence $\gamma(\mathbf{x})$ shown in the figure is sampled at integer lattice points to yield $\gamma_{\mathrm{d}}(\mathbf{x})$, the wrapped phase is $\pi$ throughout, but unwrapping across the absolute phase singularity surrounding the point at $x_{1}=x_{2}=0$ cannot pick up the $2\pi$ jump in absolute phase. Thus, it is not sufficient for the coherence and wrapped phase to change slowly. Across the absolute phase singularity, reconstruction breaks down because Itoh’s criterion is violated. The violation is analogous to a discontinuous displacement or elevation difference, but the discontinuity is always an integer multiple of $2\pi$.
@@
Certain violations of Itoh’s criterion can be detected through path dependence of phase unwrapping [27]. A nonzero unwrapped phase along a loop identifies a violation. As in the continuous case, nonzero unwrapped phases occur as integer multiples of $2\pi$. Elementary 2 $\times$ 2 pixel loops (see Fig. 9a for two dimensions) can yield nonzero $N$.
An elementary loop with nonzero $N$ is said to encircle a phase singularity, in analogy to the continuous case. In InSAR, such a loop and $N$ are referred to as residues [1, 2]. In two dimensions, a standard loop orientation can be chosen [16], assigning generic phase singularities a value of $N=1$ (positive) or $N=-1$ (negative). The elementary loop integral is a discrete exterior derivative operation such as the curl [27, 1]. In this discrete setting, the only non-generic singularity has $N=2$, an example of which is shown in Fig. 9a). Each elementary loop segment contributes $\Delta\phi_{\mathrm{d}}=\pi$. The same result would be obtained for the inverse loop because the unwrapped phase is restricted to $(-\pi,\pi]$. This violation of antisymmetry, $\pi=\Delta\phi_{\mathrm{d}}(\mathbf{x},\mathbf{x}+\Delta_{i})\neq-\Delta\phi_{\mathrm{d}}(\mathbf{x}+\Delta_{i},\mathbf{x})$, challenges discrete approximations of the exterior derivative [27, 1], but these complications are usually neglected because they do not arise generically.
If $X_{\mathrm{d}}$ contains no holes, elementary loops suffice to detect path dependence [29]. If $X_{\mathrm{d}}$ contains holes covering lattice points (e.g., masked water bodies), path dependence is detected through loops encircling the hole (Fig. 9b).
Itoh violations due to absolute phase singularities in $X_{\mathrm{d}}$ cannot be detected using path dependence. As in the continuous case (Sec. 7.1.1), unpatchable absolute phase singularities are not directly visible in $X_{\mathrm{d}}$. The discrete unwrapping of (29) is insensitive to phase jumps by integer multiples of $2\pi$ across these absolute phase singularities.
Figure 9: Loops identify path dependence in discrete, two-dimensional interferograms through a non-zero winding number $N$, printed inside a loop. a) Elementary loops are shown around two generic phase singularities with $N=\pm 1$ and a non-generic singularity with $N=2$. b) The large loop with $N=1$ encloses a hole rather than a phase singularity.
@@
Branch cuts should cover absolute phase singularities, but branch cut placement strategies are geared toward erasing path dependence, not absolute phase singularities. These strategies generally connect positive and negative phase singularities. Let us quickly review the rationale, accounting for noise. Consider a phase field meeting (31) and apply a localized, noise-like perturbation in a simply connected interior subset of $X_{\mathrm{d}}$. As an example, the pixel at $(x_{1},x_{2})=(1,2)$ in Fig. 9a) may be thought of as having been perturbed, which induced path dependence. In two dimensions, such localized path dependence implies residues whose total charge is zero because loops surrounding the perturbation remain unaffected [17]. The pair of residues with $N=\pm 1$ in Fig. 9a) has vanishing total charge. Generically, paired residues of opposite charge nucleate by bifurcation as the perturbation magnitude increases [17, 19]. Phase unwrapping errors can be contained by branch cuts connecting such a residue pair, or, more generally, residues with vanishing total charge [27]. This approach fails for unpatchable absolute phase singularities because, ultimately, absolute phase singularities generally bear no relation with residues and are invisible in the discrete wrapped phase (Fig. 8). Similar challenges exist in higher dimensions. In three dimensions, localized disturbances induce path dependence around phase singularities that are generically loop shaped [20, 19] and could be contained by a branch-cut membrane spanning the loop. In summary, absolute phase singularities should be covered by branch cuts, but the placement of branch cuts based on the observed interferogram is thus challenging across dimensions.

### v3

@@ 6.3 Loops in the punctured complex plane
$\int_{g}d\varphi_{W^{+}}=\int_{\gamma\circ g}d\varphi_{\mathbb{C}^{+}}\,.$ (27)
Evaluation is facilitated by $d\varphi_{\mathbb{C}^{+}}$ being closed [weintraub14], $\mathrm{d}d\varphi_{\mathbb{C}^{+}}=0$. (From this, the earlier claim that $d\varphi_{W^{+}}$ is closed follows, as the pullback commutes with the exterior derivative.) The one-form $d\varphi_{\mathbb{C}^{+}}$ in $\mathbb{C}^{+}$ is the canonical example of a closed one-form that is not exact, i.e., it cannot be written as the exterior derivative of a zero-form [fulton97].
The integral $\int_{\gamma\circ g}d\varphi_{\mathbb{C}^{+}}$ is determined by how often the loop $\gamma\circ g$ in $W^{+}$ winds around the origin. The one-form $d\varphi_{\mathbb{C}^{+}}$ measures the angle as seen from the origin [fulton97], yielding an integer multiple of $2\pi$ for a loop $h$ in $\mathbb{C}^{+}$, viz.
$\int_{h}d\varphi_{\mathbb{C}^{+}}=2\pi N(h)\,.$ (28)
The winding number $N(h)\in\mathbb{Z}$ counts the net number of turns $h$ makes around the origin and is positive for counterclockwise turns [fulton97]. It follows that the unwrapped phases along two paths differ by an integer multiple of $2\pi$, namely $N(\gamma\circ g)$ of the corresponding loop $g\in W^{+}$.
Two loops in $\mathbb{C}^{+}$ have the same winding number if and only if they are homotopic to each other [fulton97], with the basepoint fixed. Consequently, a loop with winding number $N$ and basepoint $1+\mathsf{j}$ can be deformed into the circular loop $h_{N}(s)\equiv\cos(2\pi Ns)+\mathsf{j}\sin(2\pi Ns)$ with $s\in[0,1]$. Thus, integrating $d\varphi_{\mathbb{C}^{+}}$ around a loop $h$ gives zero if and only if $N(h)=0$, or equivalently, if and only if $h$ can be shrunk continuously to the constant loop $h_{0}$, i.e. $h\simeq h_{0}$.
In the patchable example from Fig. 8, the two loops $\gamma\circ g$ and $\gamma\circ g^{\prime}$ are seen in the panel e) to both have a winding number of zero: neither encircles the origin, implying they are homotopic to each other and to the constant loop $h_{0}$. Conversely, the loop $\gamma\circ g$ in the unpatchable example in panel f) has a winding number of one, as it encircles the origin once in the positive direction.
@@ 6.4 Identifying patchability in $W^{+}$
The absolute phase singularities $A^{\prime}$ are thus patchable if any arbitrary loop $g\in W^{+}$ with basepoint $\mathbf{w}_{1}$ gets mapped to a loop $\gamma\circ g$ in $\mathbb{C}^{+}$ with $N(\gamma\circ g)=0$. No loop $g\in W^{+}$ can have $N(\gamma\circ g)\neq 0$. This raises two questions, namely what characterizes loops $g$ in $W^{+}$ with $N(\gamma\circ g)\neq 0$, and how can we tell whether such loops exist?
Two properties characterize a loop $g\in W^{+}$ with $N(\gamma\circ g))\neq 0$. First, it encircles a phase singularity (or other hole in $W^{+}$) such that it cannot be continuously shrunk to the constant loop $\varepsilon_{\mathbf{w}_{1}}$. If it were homotopic to $\varepsilon_{\mathbf{w}_{1}}$, the integral would evaluate to zero due to (26), which in turn implies $N(\gamma\circ g)=0$ due to (27), a contradiction. Second, it intersects $A^{\prime}$, the set of absolute phase singularities on which the phase is defined. To see this, assume it did not, so that its support is in $W^{\mathrm{a}}$. However, $d\varphi_{W^{+}}=\mathrm{d}\phi^{\mathrm{a}}$ and thus exact in $W^{\mathrm{a}}$, implying $N(\gamma\circ g)$ = 0, yielding a contradiction. Both properties apply to $g$ in Fig. 8b.
To test whether a loop $g$ with $N(\gamma\circ g)\neq 0$ exists, we turn to the fundamental group of $W^{+}$ [hatcher01]. The fundamental group $\pi_{1}(W^{+},\mathbf{w}_{1})$ of $W^{+}$ comprises all homotopy equivalence classes of loops with basepoint $\mathbf{w}_{1}$. The equivalence class represented by the loop $g$ is denoted by $[g]$, and a loop $g^{\prime}$ is in this class, $g^{\prime}\in[g]$, if $g$ and $g^{\prime}$ are homotopic, i.e. $g\simeq g^{\prime}$. The group operation is concatenation, and $\varepsilon_{\mathbf{w}_{1}}$ is the identity element. No loop $g$ with $N(\gamma\circ g)\neq 0$ exists if and only if $N(\gamma\circ h)=0$ for all $[h]$ in the fundamental group $\pi_{1}(W^{+},\mathbf{w}_{1})$. If $H$ is a set of generators in a presentation of the fundamental group, the test can be restricted to all $h]\in H$. Another way to interpret this test is to require that $\mathrm{im}(\gamma_{*})=\{h_{0}\}$, i.e., that the image of the induced homomorphism of $\gamma$, $\gamma_{*}$, only comprise $h_{0}$, the constant loop in $\mathbb{C}^{+}$. The induced homomorphism thus needs to map $\pi_{1}(W^{+},\mathbf{w}_{1})$ to the trivial group (cf. Proposition 1.33 in [hatcher01] for a covering space perspective). While identifying $\pi_{1}(W^{+},\mathbf{w}_{1})$ is generally difficult, the test is easy to apply.
@@ 7 Reconstructing the absolute phase
We turn to reconstructing the absolute phase from InSAR observations. The reconstructed absolute phase may serve as input to displacement or elevation retrievals, but the reconstruction does not address whether the absolute phase permits such geometric interpretation as a range difference.
Under what conditions can the absolute phase be reconstructed from discrete InSAR measurements? A general two-step approach [zebker98, rosen00] for a multidimensional interferogram is to i) unwrap the phase to obtain the absolute phase up to a uniform offset, an integer multiple of $2\pi$; and ii) estimate this offset using such strategies as multifrequency observations.
@@ 7.1 Multidimensional interferograms
In a multidimensional interferogram, the absolute phase may be reconstructed through the slow variation of the phase along these dimensions, most commonly space or time [ghiglia98]. In an ideal scenario, phase unwrapping reconstructs the absolute phase up to an integer multiple of $2\pi$, the estimation of which is deferred to Sec. 7.2. For now, the focus is on the conditions under which the absolute phase can be reconstructed up to a single constant by unwrapping, treating idealized continuous and then discretely sampled observations.
@@ 7.1.1 Continuous unwrapping amid singularities
The complex coherence $\gamma$ is assumed to vary smoothly in the space $X\cong S\subset\mathbb{R}^{n}$. For instance, $n=2$ for a spatial interferogram. Discontinuities in $\gamma$ (e.g., layover) have been excised from $X$. Let the set $X^{+}=X\setminus\gamma^{-1}\left[\{0\}\right]$, obtained by removing phase singularities, be open, path connected and locally path connected [hatcher01]. The wrapped phase $\varphi=\mathrm{arg}\gamma$ is assumed to arise congruently from a smooth absolute phase field $\phi^{\mathrm{a}}:X^{\mathrm{a}}\rightarrow\mathbb{R}$, where $X^{\mathrm{a}}=X\setminus A$, the space $X$ excluding all absolute phase singularities $A$. For instance, in differential InSAR with given observation times $t_{1}$ and $t_{2}$, a smooth $\phi^{\mathrm{a}}(\mathbf{x})$ over $X$ is obtained whenever the unobserved coherence $\gamma(\mathbf{x},t)$ at $t\in[t_{1},t_{2}]$ with fixed primary $t_{1}$ is a smooth function of space $\mathbf{x}$ and time $t$; however, vanishing coherence at $\mathbf{x}$ and some intermediate $t$ induces an absolute phase singularity at $\mathbf{x}$ at secondary time $t_{2}$.
@@
The phase unwrapping considered here consists of integrating $d\varphi_{X^{+}}$ along a path. The differential one-form $d\varphi_{X^{+}}$ is defined in analogy to Sec. 2.2.1 and also closed. Given an observed $\gamma(\mathbf{x})$ on $X^{+}$, an unwrapped phase difference between points $\mathbf{x}$ and $\mathbf{x}^{\prime}$ can be computed through $\Delta\phi(\Gamma)=\int_{\Gamma}d\varphi_{X^{+}}$ along a path $\Gamma:[0,1]\rightarrow X^{+}$ with $\Gamma(0)=\mathbf{x}$ and $\Gamma(1)=\mathbf{x}^{\prime}$. For fixed endpoints $\mathbf{x}$ and $\mathbf{x}^{\prime}$, $\Delta\phi$ may depend on the path.
@@
The absolute phase can be reconstructed up to a uniform offset if all unwrapping paths are restricted to $X^{\mathrm{a}}$. Reconstructability in a path-connected space $X^{\mathrm{a}}$ follows because the observed differential form $d\varphi_{X^{+}}$ is identical to $\mathrm{d}\phi^{\mathrm{a}}$ on $X^{\mathrm{a}}$ due to the assumed smoothness and congruence. The goal, then, is to avoid absolute phase singularities during unwrapping.
@@
Are absolute phase singularities related to branch cuts? Branch cuts are barriers erected during phase unwrapping to circumvent path dependence [goldstein88, rosen00]. For instance, in Fig. 9c, the two phase singularities could be connected by a branch cut, which enables recovery of an unwrapped phase field outside the branch cut. However, this unwrapped phase bears no resemblance to the absolute phase in Fig. 9d. To recover the absolute phase, branch cuts should be placed at absolute phase singularities. Doing so can be difficult, as absolute phase singularities need not have observable manifestations (Fig. 9a). However, the strategy of linking the two phase singularities with a branch cut is appropriate when path dependence is induced by localized noise, to be studied in discrete data.
@@ 7.1.2 Discrete unwrapping amid singularities
Path-based phase unwrapping proceeds along coordinate-parallel segments connecting adjacent lattice points. A space $X_{\mathrm{d}}\subset\mathbb{Z}^{n}$ is called lattice path connected if any pair of points can be connected by such a path traversing adjacent pixels. Phase unwrapping along a unit step in $x_{i}$, denoted by $\Delta_{i}$, evaluates
@@
known as Itoh’s criterion [itoh82]. Itoh’s criterion requires the absolute phase to change slowly between adjacent lattice points and can be violated due to steep gradients (undersampling) or apparent discontinuities [zebker98, pepin24]. The required slow variation warrants re-interpretation amid absolute phase singularities.
Itoh’s criterion can be violated due to absolute phase singularities, in addition to rapid phase changes or noise. An example is shown in Fig. 9a. Assuming the continuous coherence $\gamma(\mathbf{x})$ shown in the figure is sampled at integer lattice points to yield $\gamma_{\mathrm{d}}(\mathbf{x})$, the wrapped phase is $\pi$ throughout, but unwrapping across the absolute phase singularity surrounding the point at $x_{1}=x_{2}=0$ cannot pick up the $2\pi$ jump in absolute phase (Fig. 9b). Thus, it is not sufficient for the coherence and wrapped phase to change slowly. Across the absolute phase singularity, reconstruction breaks down because Itoh’s criterion is violated. The violation is analogous to a discontinuous displacement or elevation difference, but the discontinuity is always an integer multiple of $2\pi$.
@@
Certain violations of Itoh’s criterion can be detected through path dependence of phase unwrapping [goldstein88]. A nonzero unwrapped phase along a loop identifies a violation. As in the continuous case, these nonzero unwrapped phases occur as integer multiples of $2\pi$. Elementary 2 $\times$ 2 pixel loops (see Fig. 10a for two dimensions) can yield nonzero $N$.
An elementary loop with nonzero $N$ is said to encircle a phase singularity, by analogy to the continuous case. In InSAR, such a loop and $N$ are referred to as residue [bamler98, rosen00]. In two dimensions, a standard loop orientation can be chosen [berry09], assigning generic phase singularities a value of $N=1$ (positive) or N=1N=-1 (negative). The elementary loop integral is a discrete exterior derivative operation such as the curl [goldstein88, bamler98]. In this discrete setting, the only non-generic singularity has $N=2$. This non-generic singularity is shown in Fig. 10a): Each elementary loop segment contributes $\Delta\phi_{\mathrm{d}}=\pi$, due to the restriction to $(-\pi,\pi]$. The same result would be obtained for the inverse loop because the unwrapped phase is restricted to $(-\pi,\pi]$. This violation of antisymmetry, $\pi=\Delta\phi_{\mathrm{d}}(\mathbf{x},\mathbf{x}+\Delta_{i})\neq-\Delta\phi_{\mathrm{d}}(\mathbf{x}+\Delta_{i},\mathbf{x})$, challenges discrete approximations of the exterior derivative [goldstein88, bamler98]. These complications are usually neglected in practice because they do not arise generically.
If $X_{\mathrm{d}}$ contains no holes, elementary loops suffice to detect path dependence [costantini98]. If $X_{\mathrm{d}}$ contains holes covering lattice points, path dependence is detected through loops encircling these holes (Fig. 10b). Such holes are commonly encountered in practice, for instance when water bodies are masked.
Itoh violations due to absolute phase singularities in $X_{\mathrm{d}}$ cannot be detected using path dependence. As in the continuous case (Sec. 7.1.1), unpatchable absolute phase singularities are not directly visible in $X_{\mathrm{d}}$. The discrete unwrapping in (29) is insensitive to phase jumps by integer multiples of $2\pi$ across these absolute phase singularities.
Figure 10: Loops identify path dependence in discrete, two-dimensional interferograms through a non-zero winding number NN, printed inside a loop. a) Elementary loops are shown around two generic phase singularities with $N=\pm 1$ and a non-generic singularity with $N=2$. b) The large loop with $N=1$ encloses a hole rather than a phase singularity.
@@
Branch cuts should cover absolute phase singularities, but branch cut placement strategies are geared toward erasing path dependence in the interferogram. These strategies generally connect positive and negative phase singularities. Let us quickly review the rationale, accounting for noise. Consider a phase field meeting (31) and apply a localized, noise-like perturbation in a simply connected interior subset of $X_{\mathrm{d}}$. As an example, the pixel at $(x_{1},x_{2})=(1,2)$ in Fig. 10a) may be thought of as having been perturbed, which induced path dependence. In two dimensions, such localized path dependence implies residues whose total charge is zero because loops surrounding the perturbation remain unaffected [freund95]. The pair of residues with $N=\pm 1$ in Fig. 10a) has vanishing total charge. Such an oppositely charged pair arises generically as the perturbation magnitude increases: the two residues nucleate together via bifurcation and then separate [freund95, adachi07]. Phase unwrapping errors can be contained by branch cuts connecting such a residue pair, or, more generally, residues with vanishing total charge [goldstein88].
Branch cut strategies connecting residue pairs fail for unpatchable absolute phase singularities because, ultimately, absolute phase singularities generally bear no relation with residues. The are invisible in the discrete wrapped phase (Fig. 9) and do not manifest through path dependence in the interferogram. Similar challenges exist in higher dimensions. In three dimensions, localized disturbances induce path dependence around phase singularities that are generically loop shaped [huntley01, adachi07] and could be contained by a branch-cut membrane spanning the loop. In summary, absolute phase singularities should be covered by branch cuts, but the placement of branch cuts based on the observed interferogram is challenging.